OpenStudy (anonymous):

(b+1/3)(b+1/5) Multiply and simplify the answer.

OpenStudy (lalaly):

is it \[(\frac{b+1}{3})(\frac{b+1}{5})\]

OpenStudy (anonymous):

\[=b^2+1/3*b+1/5*b+1/15=b^2+8/15*b+1/15\]

OpenStudy (anonymous):

no, it's (b as the whole number + 1/3)(b as the whole number + 1/3)

OpenStudy (anonymous):

\[\frac{1}{15} \left(15 b^2+8 b+1\right) \]